OpenStudy (anonymous):

Given that y=4e -x/500 find the value of x for which y=6

OpenStudy (anonymous):

\[Y = 4e ^{-x/500} \] :|

OpenStudy (anonymous):

\[6=4e^{-\frac{x}{500}}\]divide by 4 \[1.5=e^{-\frac{x}{500}}\]write in equivalent logarithmic form \[-\frac{x}{500}=\ln(1.5)\]multiply by -500 \[x=-500\ln(1.5)\]

OpenStudy (anonymous):

Plug y value in, then isolate e Then take ln to solve for x!